OpenStudy (banjoman):

Eleven subtracted from four times a number is 25. What is the number? medal given . how do you write that as an expression? and what does X equal?

OpenStudy (brightemmerald):

11-(4x)=25

OpenStudy (brooke..help00):

X= 4(25)-11 X= 100-11 X=89 Please medal